OpenStudy (anonymous):

Would anyone be able to factor this with work? a^2 + 10ab + 25b^2 - c^2

OpenStudy (cruffo):

It looks like a difference of squares :)

OpenStudy (anonymous):

what was the formula again? (a^2-b^2) =?

OpenStudy (cruffo):

Right: a^2 - b^2 = (a+b)(a-b)

OpenStudy (anonymous):

right, so would i just pair each term? like a^2 + 10ab?

OpenStudy (cruffo):

Consider factoring the first part a^2 + 10ab + 25b^2 = (a + 5)^2

OpenStudy (anonymous):

oh thats right, sorry, im just a bit rusty

OpenStudy (cruffo):

Sorry, missed the b didn't I :) a^2 + 10ab + 25b^2 = (a + 5b)^2

OpenStudy (anonymous):

thanks for the assistance, i greatly appreciate it, both of you

OpenStudy (anonymous):

just making sure, how do i know where to add the c?

OpenStudy (anonymous):

would i just follow the formula

OpenStudy (cruffo):

@cornitodisc is using "factoring by grouping" I was thinking that by factoring the first part you could have a difference of squares: (a+5b)^2 - c^2 = (a+5b + c)(a+5b - c)

OpenStudy (anonymous):

thats the right answer according to my teacher, when i looked at it the first time however, i was trying to break it into two groups. i didnt know what to do after i saw that 25b^2 and c^2 didnt have a common factor...

OpenStudy (cruffo):

I would say that usually when you see 4 terms factoring by breaking it up into two groups would be the first thing to try. The extra letters, and the particular numbers gave me the idea to do it using another method.

OpenStudy (anonymous):

i wouldnt have been able to see that haha

OpenStudy (cruffo):

Let me see... (a^2 + 10ab ) + (25b^2 - c^2) factor an a out of the first group: a(a+10b) + (25b^2 - c^2) There is no common factor in the second group, but it does fit the pattern x^2 - y^2, so a(a+10b) + (5b + c)(5b - c)

OpenStudy (anonymous):

ah,,i see,,sorry

OpenStudy (anonymous):

its ok, ty for the input

OpenStudy (cruffo):

Yah, I don't see how to continue from here.

OpenStudy (anonymous):

lol thats what stumped me

OpenStudy (cruffo):

It's all the extra letters!!! What I saw was the "perfect square trinomial" a^2 + 10ab + 25b^2, so I knew that part factored to (something)^2. Then I saw the "(something)^2-c^2" and thought "difference of squares"...

OpenStudy (anonymous):

I get it, haha, factoring is such a pain to me, thanks for the assistance!

OpenStudy (cruffo):

Cool! Thank you!

OpenStudy (anonymous):

:D